] Zippy Company has 50,000 shares of stock outstanding and projects the following net income over...

]

Zippy Company has 50,000 shares of stock outstanding and projects the following net income over the next four years.

Year Net Income
2020 $90,000
2021 $120,000
2022 $85,000
2023 $135,000

Zippy pays a small regular dividend of $0.75 plus an extra 40% of profits when net income is higher than $100,000. What is the dividend payment for 2023?
Answer #1

Solution:

The net income is $135,000 in the year 2023 and it is more than $100,000. It means that the company will pay 40% of the total profit as dividends along with $0.75 per share.

Total share of the company = 50,000

Total Dividend in dollars = 50000 *0.75 + 135000*40% = 91,500

Dividend per share = Total dividend / Shares = 91,500 / 50,000 = $1.83